# Fluid dynamics, Statistical physics, Theoretical physics

### v1

@@ Appendix A Essentials of the loop space calculus
The field strength $F_{\mu\nu}$ is generated by the antisymmetric part of the second dot derivative. For details, see [Migdal(2025a)], Sec. 3.6, Eq. (25).
@@
The covariant derivative of the field strength is generated by the discontinuity of the dot derivative, $\partial_{\mu}(t)\equiv\frac{\delta}{\delta\dot{C}_{\mu}(t-)}-\frac{\delta}{\delta\dot{C}_{\mu}(t+)}$. For details, see [Migdal(2025a)], Sec. 3.7, Eq. (31).
@@
The dot derivative discontinuity $\partial_{\mu}(t)$ annihilates the Wilson loop. For details, see [Migdal(2025a)], Sec. 3.9, Eq. (33).
@@
The area derivative acts as a first-order differential operator, satisfying the standard Leibniz rules. For details, see [Migdal(2025a)], Sec. 3.10, Eqs. (39-41).
@@
The calculus contains a kinematical Bianchi identity. For details, see [Migdal(2025a)], Sec. 3.11, Eq. (43).
@@
The dot derivatives combine to form the loop operator $\mathcal{L}_{C}$, yielding a closed, linear diffusion equation for the Wilson loop functional. For details, see [Migdal(2025a)], Sec. 3.12, Eqs. (45-48).
@@ Appendix D The Hodge-Dual Matrix Surface
This section summarizes the definition of the Hodge-dual matrix-valued minimal surface. The full derivations and properties are detailed in Sec. 4.2 of [Migdal(2025a)].
@@
The surface is embedded in $\mathbb{R}^{4}\otimes\mathbb{R}^{4}$ with matrix-valued coordinates $X_{\mu}^{A}(\xi_{1},\xi_{2})$. For details, see [Migdal(2025a)], Sec. 4.2, Eq. (82).
@@
The area functional is defined with a Hodge-duality constraint. For details, see [Migdal(2025a)], Sec. 4.2, Eqs. (83-86).
@@
The surface boundary is mapped from the loop $C(\theta)$ using ’t Hooft projectors. For details, see [Migdal(2025a)], Sec. 4.2, Eqs. (91-97).
@@
The conformal gauge is fixed by a condition on the induced metric at the boundary. For details, see [Migdal(2025a)], Sec. 4.2, Eq. (107).
@@ Appendix E The loop equation for the minimal surface and its solution
This section summarizes how the Hodge-dual surface provides an exact, non-singular solution to the fixed-point loop equation. The full derivation is in Sec. 4.3 of [Migdal(2025a)].
@@
The area functional is defined as the minimum of a constrained action. For details, see [Migdal(2025a)], Sec. 4.3, Eqs. (111-113).
@@
The variation of the area functional $\delta|S|$ yields the area derivative. For details, see [Migdal(2025a)], Sec. 4.3, Eqs. (121, 122).
@@
The Hodge-duality of the area derivative requires the Lagrange multipliers to vanish at the boundary. For details, see [Migdal(2025a)], Sec. 4.3, Eqs. (123, 124).
@@
The area derivative of the Hodge-dual surface is, by construction, proportional to the self-dual tensor $t_{\mu\nu}$. For details, see [Migdal(2025a)], Sec. 4.3, Eq. (125).
@@
This identity, combined with the kinematical Bianchi identity (from Appendix A), proves that the loop operator annihilates the area functional. For details, see [Migdal(2025a)], Sec. 4.3, Eq. (126).
@@
Therefore, the exponential of the Hodge-dual area is an exact, non-singular solution to the fixed-point loop equation. For details, see [Migdal(2025a)], Sec. 4.3, Eq. (127).
@@
The Hodge-dual minimal surface ac be analytically found for an arbitrary non-intersecting planar loop. The main result is the following theorem.
@@
The proof, detailed in [Migdal(2025a)] (Sec. 4.7), utilizes the conformal invariance of the dual area functional to map the unit circle to the planar loop $C$. The surface is parametrized using coordinates $\xi=(x_{1},x_{2})$ inside the loop.
@@
$X^{A}_{\dot{l}}=Y^{A}_{\dot{l}}(\xi),\quad\dot{l}=3,4,\quad\xi\in\mathcal{D},\quad C=\partial\mathcal{D}$ (105)
The new field $Y^{A}_{\dot{l}}(\xi)$ satisfies trivial boundary conditions, $Y^{A}_{\dot{l}}(\xi\in C)=0$. The proof demonstrates that the area functional is minimized at $Y^{A}_{\dot{l}}(\xi)=0$.
The resulting minimal area is a universal constant multiplied by the planar area $|\mathcal{D}|$. For details, see [Migdal(2025a)], Sec. 4.7, Eqs. (166-167).
$\sqrt{\tfrac{1}{2}\Sigma_{lm}^{2}}=2\sqrt{2},$ (106)
$\left|S\left[C_{\mathrm{planar}}\right]\right|=2\sqrt{2}\,|\mathcal{D}|.$ (107)
@@
For an arbitrary loop $C$, the self-dual surface area $|S|$ is bounded by the unconstrained minimal area in the embedding space. These bounds are crucial for confinement. The full proofs are provided in [Migdal(2025a)] (Sec. 4.8).
@@
The area of the constrained matrix surface (i.e., with the duality constraint) is not smaller than that of the minimal unconstrained matrix surface with the same bounding curve. This follows from the fact that the constrained configuration space is a subset of the unconstrained space. See [Migdal(2025a)], Sec. 4.8, Eq. (168).
@@
The area of the unconstrained matrix surface is positive but not larger than four times the minimal surface area in Euclidean space ($A_{4}[C]$) with the same bounding curve $C\subset\mathbb{R}^{4}$. The proof relies on the inequality $\left|\Sigma_{\mu\nu}\right|\leq\sum_{A}\sqrt{\det\lVert g^{A}_{lm}\rVert}$. This leads to the final bound: See [Migdal(2025a)], Sec. 4.8, Eq. (172).
@@
A key conceptual issue, first pointed out by E. Witten, is that a single Hodge-dual surface with a fixed chirality ($\lambda=\pm 1$) would violate parity. The full analysis is in [Migdal(2025a)] (Sec. 4.9).
To restore parity and preserve the crucial factorization property for separated loops, the solution must be the exponential of the sum of both the self-dual and anti-self-dual areas. See [Migdal(2025a)], Sec. 4.9, Eq. (176).

### v2

@@ 4. The Closed Loop Condition
The field strength $F_{\mu\nu}$ is generated by the antisymmetric part of the second dot derivative. For details, see [Migdal(2025b)], Sec. 3.6, Eq. (25).
@@
The covariant derivative of the field strength is generated by the discontinuity of the dot derivative, $\partial_{\mu}(t)\equiv\frac{\delta}{\delta\dot{C}_{\mu}(t-)}-\frac{\delta}{\delta\dot{C}_{\mu}(t+)}$. For details, see [Migdal(2025b)], Sec. 3.7, Eq. (31).
@@
The dot derivative discontinuity $\partial_{\mu}(t)$ annihilates the Wilson loop. For details, see [Migdal(2025b)], Sec. 3.9, Eq. (33).
@@
The area derivative acts as a first-order differential operator, satisfying the standard Leibniz rules. For details, see [Migdal(2025b)], Sec. 3.10, Eqs. (39-41).
@@
The calculus contains a kinematical Bianchi identity. For details, see [Migdal(2025b)], Sec. 3.11, Eq. (43).
@@
The dot derivatives combine to form the loop operator $\mathcal{L}_{C}$, yielding a closed, linear diffusion equation for the Wilson loop functional. For details, see [Migdal(2025b)], Sec. 3.12, Eqs. (45-48).
@@ Appendix D The Hodge-Dual Matrix Surface
This section summarizes the definition of the Hodge-dual matrix-valued minimal surface. The full derivations and properties are detailed in Sec. 4.2 of [Migdal(2025b)].
@@
The surface is embedded in $\mathbb{R}^{4}\otimes\mathbb{R}^{4}$ with matrix-valued coordinates $X_{\mu}^{A}(\xi_{1},\xi_{2})$. For details, see [Migdal(2025b)], Sec. 4.2, Eq. (82).
@@
The area functional is defined with a Hodge-duality constraint. For details, see [Migdal(2025b)], Sec. 4.2, Eqs. (83-86).
@@
The surface boundary is mapped from the loop $C(\theta)$ using ’t Hooft projectors. For details, see [Migdal(2025b)], Sec. 4.2, Eqs. (91-97).
@@
The conformal gauge is fixed by a condition on the induced metric at the boundary. For details, see [Migdal(2025b)], Sec. 4.2, Eq. (107).
@@ Appendix E The loop equation for the minimal surface and its solution
This section summarizes how the Hodge-dual surface provides an exact, non-singular solution to the fixed-point loop equation. The full derivation is in Sec. 4.3 of [Migdal(2025b)].
@@
The area functional is defined as the minimum of a constrained action. For details, see [Migdal(2025b)], Sec. 4.3, Eqs. (111-113).
@@
The variation of the area functional $\delta|S|$ yields the area derivative. For details, see [Migdal(2025b)], Sec. 4.3, Eqs. (121, 122).
@@
The Hodge-duality of the area derivative requires the Lagrange multipliers to vanish at the boundary. For details, see [Migdal(2025b)], Sec. 4.3, Eqs. (123, 124).
@@
The area derivative of the Hodge-dual surface is, by construction, proportional to the self-dual tensor $t_{\mu\nu}$. For details, see [Migdal(2025b)], Sec. 4.3, Eq. (125).
@@
This identity, combined with the kinematical Bianchi identity (from Appendix A), proves that the loop operator annihilates the area functional. For details, see [Migdal(2025b)], Sec. 4.3, Eq. (126).
@@
Therefore, the exponential of the Hodge-dual area is an exact, non-singular solution to the fixed-point loop equation. For details, see [Migdal(2025b)], Sec. 4.3, Eq. (127).
@@
The Hodge-dual minimal surface can be analytically found for an arbitrary non-intersecting planar loop. The main result is the following theorem. We present this theorem here, correcting the proof in [Migdal(2025b)], where there was a typo in the boundary conditions for $Y$.
@@
Proof.
Let us use the conformal invariance of the dual area functional (101) and map the unit circle onto the given planar bounding loop $\mathbb{S}_{1}\mapsto C$. After that, the minimal surface can use the coordinates $\xi=(x_{1},x_{2})$ inside that planar loop $C$ as internal coordinates of the surface. In the context of our surface embedded into $\mathbb{R}^{4}\otimes\mathbb{R}^{4}$, this parametrization reduces to
@@
$X^{A}_{\dot{l}}=Y^{A}_{\dot{l}}(\xi),\quad\dot{l}=3,4,$ (121)
$\xi\in\mathcal{D},\quad C=\partial\mathcal{D},$ (122)
where $\Lambda^{A}_{\mu\nu}$ are defined in (105), and the new field $Y^{A}_{\dot{l}}(\xi)$ satisfies the boundary conditions:
[…]
The area functional is
$|S|=\int_{\mathcal{D}}d^{2}\xi\sqrt{\tfrac{1}{2}\Sigma_{lm}^{2}+\Sigma_{l\dot{l}}^{2}+\tfrac{1}{2}\Sigma_{\dot{l}\dot{m}}^{2}},$ (124)
$\Sigma_{lm}=e_{ab}\Lambda^{A}_{al}\Lambda^{A}_{bm},$ (125)
$\Sigma_{l\dot{l}}=e_{ab}\Lambda^{A}_{al}\,\partial_{b}Y^{A}_{\dot{l}},$ (126)
$\Sigma_{\dot{l}\dot{m}}=e_{ab}\,\partial_{a}Y^{A}_{\dot{l}}\,\partial_{b}Y^{A}_{\dot{m}}.$ (127)
With the Dirichlet boundary values (123) the constrained (self–dual) minimizer of the area functional (124) is the linear map
$X^{A}_{\mu}(\xi)=\Lambda^{A}_{\mu m}\,\xi_{m},\qquad\xi\in D.$ (128)
This embedding automatically satisfies both the Hodge–duality constraint and the corrected boundary conditions. Its area tensor is constant,
$\Sigma_{\mu\nu}=\sum_{A}\Lambda^{A}E(\Lambda^{A})^{\!T}=2\!\begin{pmatrix}0&1&0&0\\[-3.0pt] -1&0&0&0\\[-3.0pt] 0&0&0&1\\[-3.0pt] 0&0&-1&0\end{pmatrix}$, (129)
which is self–dual ($\Sigma=\Sigma$) and yields $\Sigma_{l\dot{l}}=0$. The Euler–Lagrange system is identically satisfied for linear $X^{A}_{\mu}(\xi)$. The corresponding area density is $\sqrt{\tfrac{1}{2}\Sigma_{\mu\nu}\Sigma_{\mu\nu}}=2\sqrt{2},$ so that
$|S|=\!\!\int_{D}\!d^{2}\xi\,2\sqrt{2}=2\sqrt{2}\,|D|.$ (130)
Hence the minimal dual area for any non–intersecting planar loop equals $2\sqrt{2}$ times the Euclidean area enclosed by the loop, and the linear map (128) provides the unique regular self–dual minimizer consistent with the boundary data. ∎
@@
For an arbitrary loop $C$, the self-dual surface area $|S|$ is bounded by the unconstrained minimal area in the embedding space. These bounds are crucial for confinement. The full proofs are provided in [Migdal(2025b)] (Sec. 4.8).
@@
The area of the constrained matrix surface (i.e., with the duality constraint) is not smaller than that of the minimal unconstrained matrix surface with the same bounding curve. This follows from the fact that the constrained configuration space is a subset of the unconstrained space. See [Migdal(2025b)], Sec. 4.8, Eq. (168).
@@
The area of the unconstrained matrix surface is positive but not larger than four times the minimal surface area in Euclidean space ($A_{4}[C]$) with the same bounding curve $C\subset\mathbb{R}^{4}$. The proof relies on the inequality $\left|\Sigma_{\mu\nu}\right|\leq\sum_{A}\sqrt{\det\lVert g^{A}_{lm}\rVert}$. This leads to the final bound: See [Migdal(2025b)], Sec. 4.8, Eq. (172).
@@
A key conceptual issue, first pointed out by E. Witten, is that a single Hodge-dual surface with a fixed chirality ($\lambda=\pm 1$) would violate parity. The full analysis is in [Migdal(2025b)] (Sec. 4.9).
To restore parity and preserve the crucial factorization property for separated loops, the solution must be the exponential of the sum of both the self-dual and anti-self-dual areas. See [Migdal(2025b)], Sec. 4.9, Eq. (176).